A hexagon has 2 sides of length x inches each, 2 sides of length 3x inches each, one side of length 4x I nches and one side of length 6 inches. If the perimeter of the hexagon is 54 inches what is the value of x?
A) 4
B)4.5
C)8
D)9

Answer by Theo(13342)

2 sides of x equals 2x.
2 sides of 3x equals 6x.
1 side of 4x equals 4x
1 side of 6 equals 6

perimeter = 54, therefore 2x + 6x + 4x + 6 = 54.
combine like terms to get 12x + 6 = 54
subtract 6 from both sides of that equation to get 12x = 48
divide both sides of that equation by 12 to get x = 48/12 = 4

the value of x = 4.

2 sides of x becomes 2x = 8
2 sides of 3x becomes 6x = 24
1 side of 4x becomes 4x = 16
1 side of 6 = 6
8 + 24 + 16 + 6 = 54

solution is confirmed to be good.

solution is selection A.